# The sup-norm problem for newforms of large level on $\operatorname{PGL}(n)$

### v1

@@ 3.2. Amplifiers
Let $\mathcal{M}_{n}(\mathbb{Z},N)$ be the set of integral matrices with last row congruent to $(0,\ldots,0,*)$ modulo $N$. For $(m,N)=1$ let
$H(m,N):=\{\gamma\in\mathcal{M}_{n}(\mathbb{Z},N)\mid\det\gamma=m\}$
and
@@
where the implicit constant depends on $n$, dependence which we suppress throughout the arguments.
@@
We choose the archimedean test function $f_{\mu}:C_{c}^{\infty}(K\backslash G/K)\longrightarrow\mathbb{C}$ defined in [BM15, Section 3]. It has compact support and is bounded $f_{\mu}\ll_{\mu,n}1$ in terms of $\mu$, where the dependence on $\mu$ is continuous. [^1] In fact, there is an explicit bound for the function $f_{\mu}$. However, it is only useful in the spectral aspect. For our purposes, we may simply bound $f_{\mu}$ by a constant depending on $\mu$, but independent of the level. Its spherical transform $\tilde{f}_{\mu}$ satisfies
$\tilde{f}_{\mu}(\mu)\geq 1$
and is non-negative on all possible spectral parameters occurring in the decomposition of $L^{2}(\Gamma_{0}(N)\backslash\mathbb{H})$. Finally, when writing $f_{\mu}(g)$ for $g\in\operatorname{GL}_{n}(\mathbb{R})$, where $\det(g)>0$, we mean $f_{\mu}(g/\det(g)^{1/n})$, thus extending the domain of $f_{\mu}$ by postulating its invariance under scalars. Now consider
[…]
@@
[…]
@@
[…]
@@
using the triangle inequality. Since $K$ is compact, we can simplify $k+O(1)$ to $O(1)$, where the implicit constant depends on $n$. We now write the compositions $T(p)\cdot T^{\prime}(q)$ as linear combinations of Hecke operators $T_{g}$. Let $z=w$ and assume that $p\neq q$. Recalling that $T(p)\cdot T^{\prime}(q)$ is the Hecke operator corresponding to
@@
[…]
@@
For unconditional bounds, one may work with the amplifier given in [BM15, (6.2)]. It uses Hecke operators attached to higher powers of primes for providing an alternative to Hypothesis (1.1). In fact, we give the slightly more precise version of this amplifier by including information on the determinantal divisors. Recall that the $j$-th determinantal divisor $\Delta_{j}(\gamma)$ of an integral matrix $\gamma$ is equal to the greatest common divisor of all $j\times j$ minors.
@@
Blomer and Maga only preserve the condition on $\Delta_{1}$ and $\Delta_{2}$ (see their definition of $S(m,l)$). These and the additional ones in the proposition above follow directly using the crucial property of the determinantal divisors, namely their invariance under right or left multiplication by elements of $\operatorname{SL}_{n}(\mathbb{Z})$ (see e.g. [New72, Thm. II.8]). Except for the proof of Proposition 6.10, these conditions are not used and we mostly consider the larger set $H(z,m,N)$ for simplicity of notation.
@@ 4.1. The normaliser of the Hecke congruence subgroup
In the theory of automorphic forms on $\operatorname{SL}_{2}(\mathbb{R})$, an Atkin-Lehner operator $S$ is an involution on space of left-$\Gamma_{0}(N)$ invariant functions. It is obtained by setting $Sf(z)=f(gz)$ for all $z\in\mathbb{H}$, where $g$ lies in the normaliser of $\Gamma_{0}(N)$ inside $\operatorname{SL}_{2}(\mathbb{R})$. This is a natural method of producing automorphisms, since the invariance of $f(z)$ under a group $\Gamma$ is equivalent to the invariance of $f(gz)$ under $g^{-1}\Gamma g$. The normaliser has been computed by Atkin and Lehner in [AL70] and an example of a non-trivial normalising element is
@@
which induces the so-called Fricke involution. In fact, the normaliser gives all automorphism of the modular curve $X_{2}(N)$, in more standard notation $X_{0}(N)$, for all $N$ up to finitely many exceptions (see [KM88]).
@@
Let $L=g\cdot L_{1}\in\mathcal{L}$, where $g\in\operatorname{GL}_{3}^{+}(\mathbb{Q})$, and assume that $\Gamma_{0}(N)$ fixes $L$. Then $g^{-1}\Gamma_{0}(N)g$ fixes $L_{1}$, so we must have $g^{-1}\Gamma_{0}(N)g\subset\operatorname{SL}_{3}(\mathbb{Z})$. Scaling $g$ by a positive rational number, we may assume that $g\in\mathcal{M}_{3\times 3}(\mathbb{Z})$. Let then $H$ be the Hermite normal form of $g$, so that
@@
for all $M\mid N$. By the definition of $L_{M}$ and using the fact that $\operatorname{Stab}(L_{1})=\operatorname{SL}_{3}(\mathbb{Z})$, we can deduce that
$q_{M}^{-1}\begin{pmatrix}1&&\\ &1&\\ &&f(M)^{-1}\end{pmatrix}\cdot g\cdot\begin{pmatrix}1&&\\ &1&\\ &&M\end{pmatrix}\in\operatorname{SL}_{3}(\mathbb{Z}),$ (4.1)
for all $M|N$. Rescaling $g$ by $q_{1}\in\mathbb{Q}$ we may assume that $q_{1}=1$. Taking $M=1$ in (4.1) and applying determinants, we deduce that $\det(g)=f(1)$. Applying determinants to all other equations, we find that
@@
In particular, for $M=N$, we have $q_{N}^{3}f(N)=Nf(1)$. Since $f(N)\mid N$, we must have $q_{N}\in\mathbb{Z}$. Let us make (4.1) more explicit. Taking $M=1$, we have
@@
Using the properties of the determinant and that $\ast$ denotes integers, we deduce that $q_{N}^{2}|\det(g)=f(1)$. Let $f(1)=q_{N}^{2}k$ for some $k\in\mathbb{Z}$. Now the last row of $g$ is divisible by $q_{N}^{2}k$ and the first two columns are divisible by $q_{N}$. By the same method we infer that $q_{N}k\cdot q_{N}\cdot q_{N}=q_{N}^{3}k$ divides $\det(g)=f(1)=q_{N}^{2}k$. Therefore $q_{N}=1$, which implies that $f(N)=Nf(1)$. Since $f(N)\mid N$, it follows that $f(1)=1$ and $f(N)=N$. Putting everything together, it follows that $g\in\Gamma_{0}(N)$. ∎
@@
We use the results of [Bor66], which imply that the normaliser of $\Gamma_{0}(N)$, being commensurable with the arithmetic group $\operatorname{PGL}_{n}(\mathbb{Z})$, lies in $\operatorname{PGL}_{n}(\mathbb{Q})$. ∎
@@ 4.2. A different perspective
An important automorphism of matrices in $\operatorname{SL}_{n}(\mathbb{R})$ is the map $g\mapsto g^{-T}$, sending a matrix to its inverse transpose. As already noted in the present paper, this map sends a lattice $L_{g}$ to its dual, but is also used to define the dual form of an automorphic form for $\operatorname{SL}_{n}(\mathbb{Z})$ (see section 9.2 in [Gol06]) or also the contragredient representation of a $\operatorname{GL}(n)$ automorphic representation.
@@
$wg^{-T}w^{-1}=-\frac{1}{\det(g)}g.$ (4.2)
In particular, the map $z\mapsto z^{-T}$ induces the identity on $\operatorname{PGL}_{2}(\mathbb{Z})\backslash\operatorname{PGL}_{2}(\mathbb{R})/\operatorname{PO}(2)$.
@@
We first prove that $T_{g}W_{N}=W_{N}T_{g}^{*}$. By a variant of the Smith normal form, we may assume that $g$ is diagonal and by a variant of the transposition anti-automorphism for $\Gamma_{0}(N)$ (generalising Lemma 4.5.2 and Theorem 4.5.3 in [Miy89], we may assume that there are matrices $\alpha_{i}$, $i=1,\ldots,k$, for some $k$, such that
@@
where $\beta_{i}=A_{N}\alpha_{i}^{-T}A_{N}^{-1}$. The proof is finished by showing that $\bigcup_{i}\Gamma_{0}(N)\beta_{i}=\Gamma_{0}(N)g^{-1}\Gamma_{0}(N)$, since this double coset corresponds to $T_{g}^{\ast}$ (s. [Gol06, Thm. 9.6.3]. Indeed,
@@
Here we made use of fundamental property (4.3) of $A_{N}$ and of the fact that $g$ is diagonal, thus commuting with $A_{N}$. Next, we prove that $W_{N}$ is self-adjoint. This can easily be seen by using a known fact about the dual forms for $\operatorname{SL}_{n}(\mathbb{Z})$. Namely, the map $f(z)\mapsto f(wz^{-T}w^{-1})$, where $w$ is the long Weyl element, is self-adjoint (one can compute directly in explicit coordinates given in [Gol06], Proposition 9.2.1 or Proposition 6.3.1). We can interpret the Fricke involution as
@@
where $m=A_{N}w^{-1}$, that is, as the composition of the dualising map above with the left-action of $m$. Since the measure on $\mathbb{H}^{n}$ is $\operatorname{GL}_{n}(\mathbb{R})$-invariant, we can make the same explicit computations and change of coordinates as for the dualising map. Since $A_{N}$ is diagonal, we easily deduce the conclusion $W_{N}^{\ast}=W_{N}$. Moreover, this interpretation of the Fricke involution and [BHM20, (45)] also prove the statement about the spectral parameters of $\phi$. To prove cuspidality it is best to work adelically, though this can be reduced again to noting the relation between $W_{N}$ and the dualising map. Namely, the form $W_{N}\phi$ generates the contragredient of the representation generated by $\phi$, which is known to be cuspidal (see e.g. [Bum97, Prop. 3.3.4]). From this perspective, it is also easy to see that $W_{N}\phi$ is a newform. In the interest of brevity, we leave out the details of adelisation. ∎
@@
As explained in [McD78], all automorphisms in the case $n>2$ are constructed out of inner automorphisms, radial automorphisms, and the inverse-transpose automorphism. Inner automorphisms cannot contribute, since we have proved that the normaliser of $\Gamma_{0}(N)$ is trivial; radial automorphisms are trivial in our context, since we consider only automorphic forms that are invariant under the centre of $\operatorname{GL}_{n}(\mathbb{R})$; and the inverse-transpose automorphism is precisely the basis for the definition given in this note.
4.3. Uniqueness of the Fricke involution
The theory of Atkin-Lehner operators for $\Gamma_{0}(N)$ shows some weaknesses already in the well-understood case $n=2$. Indeed, one can only define Atkin-Lehner operators for divisors $M$ of the level $N$, such that $M$ and $N/M$ are coprime. More precisely, there are no operators induced by matrices with determinant equal $M\mid N$, such that $(M,N/M)\neq 1$ (see [AL70, p. 138]).
This phenomenon creates difficulties in applications when considering powerful levels, as already noted in the historical context of the sup-norm problem. In the present section, we see that these difficulties only get more problematic in higher rank (see Remark 4.5). In fact, the only Atkin-Lehner operator for $n>2$, according to our definition, is the Fricke involution.
Proposition 4.4.
@@ 4.3. Uniqueness of the Fricke involution
We apply the same ideas as in the proof of Theorem 4. Again the proof is done for $n=3$, merely for simplicity and clarity. One can check that $\Gamma_{0}(N)^{T}$ stabilises the lattices
$L_{M^{-1}}=\langle e_{1},e_{2},M^{-1}e_{3}\rangle=\operatorname{diag}(1,1,M^{-1})L_{1}$
for all divisors $M\mid N$. It follows that $\Gamma_{0}(N)$ must stabilise (up to scalars) the lattices $gL_{M^{-1}}$. By Lemma 4.1 determining the fixed points of $\Gamma_{0}(N)$, we have
$gL_{M^{-1}}=q_{M}L_{f(M)},$
with $f(M)\mid N$. We normalise $g$ by a rational number so that $q_{1}=1$. The equations above imply that
$g\in q_{M}\operatorname{diag}(1,1,f(M))\operatorname{SL}_{3}(\mathbb{Z})\operatorname{diag}(1,1,M),$ (4.4)
using that the stabiliser of $L_{1}$ is $\operatorname{SL}_{3}(\mathbb{Z})$. Let us take determinants and deduce that
$\det g=q_{M}^{3}\cdot f(M)\cdot M.$ (4.5)
By our assumption, $\det g=f(1)$. Take $M=N$ in (4.5) and note that
$q_{N}^{-3}=\frac{f(N)N}{f(1)}.$
Since $f(1)\mid N$, we deduce that $q_{N}^{-3}\in\mathbb{Z}$, so $d:=q_{N}^{-1}\in\mathbb{Z}$. Using this notation we have $d^{3}f(1)=f(N)N$. Now we use the matrix equation for $M=1$ and $M=N$ to find that
$g=\begin{pmatrix}&&\\ &&\\ f(1)\ast&f(1)\ast&f(1)\ast\end{pmatrix}\quad\text{and}\quad g=\begin{pmatrix}&&\frac{N}{d}\ast\\ &&\frac{N}{d}\ast\\ \frac{f(N)}{d}\ast&\frac{f(N)}{d}\ast&\frac{f(N)N}{d}\ast,\end{pmatrix}$ (4.6)
where $\ast$ stands for unknown integers and the rest of the matrices are also filled by integers. We claim that
$d\mid N.$
[…]
$g=\begin{pmatrix}&&p\ast\\ &&p\ast\\ f(1)\ast&f(1)\ast&f(1)p^{2}\ast\end{pmatrix}.$
It would follow that $f(1)\cdot p\mid\det(g)=f(1)$, but this is a contradiction. Therefore $d=1$. We infer that $f(1)=Nf(N)$, so considering divisibility we must have $f(1)=N$ and $f(N)=1$. This implies that $\det g=N$ and that the last row and column of $g$ are divisible by $N$. Thus $g$ is of the form
$g=\begin{pmatrix}\alpha_{1}&\alpha_{2}&N\alpha_{3}\\ \beta_{1}&\beta_{2}&N\beta_{3}\\ N\gamma_{1}&N\gamma_{2}&N\gamma_{3}\end{pmatrix}$
with $\alpha_{i},\beta_{i},\gamma_{i}\in\mathbb{Z}$. Since $\det(g)=N$, it must be that $\gamma_{3}$ is coprime to $N$ and that $(\alpha_{3},\beta_{3},\gamma_{3})=1$. In fact, put these together to have $(N\alpha_{3},N\beta_{3},\gamma_{3})=1$. Now take $x,y,z\in\mathbb{Z}$ such that
$xN\alpha_{3}+yN\beta_{3}+z\gamma_{3}=1.$
Then $(xN,yN,z)=1$, so we can find a matrix $u\in\Gamma_{0}(N)$ with last row equal to $(xN,yN,z)$. It follows from the above that the entry in the lower right corner of $u\cdot g$ is equal to $N$. By doing row manipulations we can find $u^{\prime}\in\Gamma_{0}(N)$ such that
$u^{\prime}g=\begin{pmatrix}\ast&\ast&0\\ \ast&\ast&0\\ N\ast&N\ast&N\end{pmatrix}.$
In this form, it is obvious that we can find another $u^{\prime\prime}\in\Gamma_{0}(N)$ so that $u^{\prime\prime}g=\operatorname{diag}(1,1,N)$. ∎
Remark 4.5.
Let us note what changes in the proof in the case $n=2$ and how this leads to the lack of Atkin-Lehner operators for powerful levels. In the notation above, we would have the equation $d^{2}f(1)=f(N)N$, where the exponent of $d$ is equal to $n$ in general. We can still prove that $d\mid N$, yet the next paragraph in the proof differs slightly. We suppose $p$ is a prime dividing $d$ such that $p^{k}\mathrel{\|}d$ is the maximal power of $p$ dividing $d$, with $k\geq 1$. As in the proof above, we deduce that $p^{2k}\mid f(N)N$ and $p^{k}\mid N$. If we were to continue the proof as above and deduce that $d=1$, we would need the step showing that $p$ divides $N/d$. This is not true in this case any more. For example, if $N$ is square free, then $k\leq 1$ and the claim in the step may not hold for certain choices of $f(N)$. In fact, solving the matrix equations eventually leads to the matrices found by Atkin and Lehner (after suitably multiplying by the long Weyl element). If $N$ is powerful, then we could have that a higher power of $p$ divides $N$. For certain choices of $d$, we can indeed deduce that $p\mid N/d$ and produce a contradiction. These choices of $d$ correspond to divisors $M$ of $N$, such that $(M,N/M)\neq 1$. Indeed, suppose that $\det(g)=f(1)=:M$, $p\mid M$ and $p\mid N/M$. Then $p$ divides $d=f(N)N/M$. If $p^{k}\mathrel{\|}d$, then applying the $p$-adic valuation to $d^{2}M=f(N)N$ and recalling that $f(N)\mid N$ shows that $p\mid N/d$. We proceed as in the proof above and derive a contradiction. This shows that there are no Atkin-Lehner operators for such divisors $M$ as above.

### v2

@@ 3.2. Amplifiers
[…]
We shall consider matrices $g\in G$ in the support of $f_{\mu}$. Such an assumption implies that $g=k+O_{n}(1)$ for some $k\in K$, where $O_{n}(1)$ refers to a matrix with entries bounded by some constant depending on $n$. Now, let $\mathcal{M}_{n}(\mathbb{Z},N)$ be the set of integral matrices with last row congruent to $(0,\ldots,0,*)$ modulo $N$. For $(m,N)=1$, let
@@
where $z^{-1}\gamma z=O(m^{1/n})$ means that $m^{-1/n}z^{-1}\gamma z$ lies in the support of $f_{\mu}$, as above. The implicit constant thus depends on $n$ and the support of $f_{\mu}$, dependence which we suppress throughout the paper.
@@
In preparation for the proof, recall that, by the work of Blomer and Maga, the spherical transform $\tilde{f}_{\mu}$ of $f_{\mu}$ satisfies $\tilde{f}_{\mu}(\mu)\geq 1$ and is non-negative on all possible spectral parameters occurring in the decomposition of $\)L^{2}(\Gamma_{0}(N)\backslash\mathbb{H}$). Additionally, when writing $f_{\mu}(g)$ for $g\in\operatorname{GL}_{n}(\mathbb{R})$, where $\det(g)>0$, we mean $f_{\mu}(g/\det(g)^{1/n})$, thus extending the domain of $f_{\mu}$ by postulating its invariance under scalars.
Now consider
$\int A_{\varpi}\cdot\tilde{f}_{\mu}(\mu_{\varpi})\phi_{\varpi}(z)\overline{\phi_{\varpi}(w)}\,d\varpi,$
@@
$\frac{1}{(pq)^{(n-1)/2}}\int\lambda(p,\varpi)x_{p}\cdot\overline{\lambda(q,\varpi)x_{q}}\cdot\tilde{f}_{\mu}(\mu_{\varpi})\phi_{\varpi}(z)\overline{\phi_{\varpi}(w)}\,d\varpi,$
@@
$S_{p,q}=\frac{x_{p}\overline{x_{q}}}{(pq)^{(n-1)/2}}\cdot T(p)T^{\prime}(q)\cdot\int\tilde{f}_{\mu}(\mu_{\varpi})\phi_{\varpi}(z)\overline{\phi_{\varpi}(w)}\,d\varpi,$
@@
using the triangle inequality. Since $K$ is compact, we can simplify $k+O(1)$ to $O(1)$, where the implicit constant depends on $n$.
We now write the compositions $T(p)\cdot T^{\prime}(q)$ as linear combinations of Hecke operators $T_{g}$. Let $z=w$ and assume that $p\neq q$. Recalling that $T(p)\cdot T^{\prime}(q)$ is the Hecke operator corresponding to
@@
$A_{\varpi_{0}}|\phi(z)|^{2}\leq\int A_{\varpi}\cdot\tilde{f}_{\mu}(\mu_{\varpi})|\phi_{\varpi}(z)|^{2}d\varpi.$
@@
For unconditional bounds, one may work with the amplifier given in [7, (6.2)]. It uses Hecke operators attached to higher powers of primes for providing an alternative to Hypothesis (1.3). In fact, we give the slightly more precise version of this amplifier by including information on the determinantal divisors. Recall that the $j$-th determinantal divisor $\Delta_{j}(\gamma)$ of an integral matrix $\gamma$ is equal to the greatest common divisor of all $j\times j$ minors.
@@
Blomer and Maga only preserve the condition on $\Delta_{1}$ and $\Delta_{2}$ (see their definition of $S(m,l)$). These and the additional ones in the proposition above follow directly using the crucial property of the determinantal divisors, namely their invariance under right or left multiplication by elements of $\operatorname{SL}_{n}(\mathbb{Z})$ (see e.g. [35, Thm. II.8]). Except for the proof of Proposition 6.10, these conditions are not used, and we mostly consider the larger set $H(z,m,N)$ for simplicity of notation.
@@ 4.1. The normaliser of the Hecke congruence subgroup
In the theory of automorphic forms on $\operatorname{SL}_{2}(\mathbb{R})$, an Atkin-Lehner operator $S$ is an involution on space of left-$\Gamma_{0}(N)$ invariant functions. It is obtained by setting $Sf(z)=f(gz)$ for all $z\in\mathbb{H}$, where $g$ lies in the normaliser of $\Gamma_{0}(N)$ inside $\operatorname{SL}_{2}(\mathbb{R})$. This is a natural method of producing automorphisms, since the invariance of $f(z)$ under a group $\Gamma$ is equivalent to the invariance of $f(gz)$ under $g^{-1}\Gamma g$. The normaliser has been computed by Atkin and Lehner in [2] and an example of a non-trivial normalising element is
@@
which induces the so-called Fricke involution. In fact, the normaliser gives all automorphism of the modular curve $X_{2}(N)$, in more standard notation $X_{0}(N)$, for all $N$ up to finitely many exceptions (see [28]).
@@
Let $L=g\cdot L_{1}\in\mathcal{L}$, where $g\in\operatorname{GL}_{3}^{+}(\mathbb{Q})$, and assume that $\Gamma_{0}(N)$ fixes $L$. Then $g^{-1}\Gamma_{0}(N)g$ fixes $L_{1}$, so we must have $g^{-1}\Gamma_{0}(N)g\subset\operatorname{SL}_{3}(\mathbb{Z})$.
Scaling $g$ by a positive rational number, we may assume that $g\in\mathcal{M}_{3\times 3}(\mathbb{Z})$. Let then $H$ be the Hermite normal form of $g$, so that
@@
for all $M\mid N$.
By the definition of $L_{M}$ and using the fact that $\operatorname{Stab}(L_{1})=\operatorname{SL}_{3}(\mathbb{Z})$, we can deduce that
(4.1) $q_{M}^{-1}\begin{pmatrix}1&&\\ &1&\\ &&f(M)^{-1}\end{pmatrix}\cdot g\cdot\begin{pmatrix}1&&\\ &1&\\ &&M\end{pmatrix}\in\operatorname{SL}_{3}(\mathbb{Z}),$
for all $M|N$.
Rescaling $g$ by $q_{1}\in\mathbb{Q}$ we may assume that $q_{1}=1$. Taking $M=1$ in (4.1) and applying determinants, we deduce that $\det(g)=f(1)$. Applying determinants to all other equations, we find that
@@
In particular, for $M=N$, we have $q_{N}^{3}f(N)=Nf(1)$. Since $f(N)\mid N$, we must have $q_{N}\in\mathbb{Z}$.
Let us make (4.1) more explicit. Taking $M=1$, we have
@@
Using the properties of the determinant and that $\ast$ denotes integers, we deduce that $q_{N}^{2}|\det(g)=f(1)$.
Let $f(1)=q_{N}^{2}k$ for some $k\in\mathbb{Z}$. Now the last row of $g$ is divisible by $q_{N}^{2}k$ and the first two columns are divisible by $q_{N}$. By the same method we infer that $q_{N}k\cdot q_{N}\cdot q_{N}=q_{N}^{3}k$ divides $\det(g)=f(1)=q_{N}^{2}k$. Therefore $q_{N}=1$, which implies that $f(N)=Nf(1)$. Since $f(N)\mid N$, it follows that $f(1)=1$ and $f(N)=N$. Putting everything together, it follows that $g\in\Gamma_{0}(N)$. ∎
@@
We use the results of [10], which imply that the normaliser of $\Gamma_{0}(N)$, being commensurable with the arithmetic group $\operatorname{PGL}_{n}(\mathbb{Z})$, lies in $\operatorname{PGL}_{n}(\mathbb{Q})$. ∎
@@ 4.2. A different perspective
An important automorphism of matrices in $\operatorname{SL}_{n}(\mathbb{R})$ is the map $g\mapsto g^{-T}$, sending a matrix to its inverse transpose. As already noted in the present paper, this map sends a lattice $L_{g}$ to its dual, but is also used to define the dual form of an automorphic form for $\operatorname{SL}_{n}(\mathbb{Z})$ (see section 9.2 in [20]) or also the contragredient representation of a $\operatorname{GL}(n)$ automorphic representation.
@@
(4.2) $wg^{-T}w^{-1}=-\frac{1}{\det(g)}g.$
[…]
@@
We first prove that $T_{g}W_{N}=W_{N}T_{g}^{*}$. By a variant of the Smith normal form, we may assume that $g$ is diagonal and by a variant of the transposition anti-automorphism for $\Gamma_{0}(N)$ (generalising Lemma 4.5.2 and Theorem 4.5.3 in [34]), we may assume that there are matrices $\alpha_{i}$, $i=1,\ldots,k$, for some $k$, such that
@@
where $\beta_{i}=A_{N}\alpha_{i}^{-T}A_{N}^{-1}$. The proof is finished by showing that $\bigcup_{i}\Gamma_{0}(N)\beta_{i}=\Gamma_{0}(N)g^{-1}\Gamma_{0}(N)$, since this double coset corresponds to $T_{g}^{\ast}$ (s. [20, Thm. 9.6.3]). Indeed,
@@
Here we made use of fundamental property (4.3) of $A_{N}$ and of the fact that $g$ is diagonal, thus commuting with $A_{N}$.
Next, we prove that $W_{N}$ is self-adjoint. This can easily be seen by using a known fact about the dual forms for $\operatorname{SL}_{n}(\mathbb{Z})$. Namely, the map $f(z)\mapsto f(wz^{-T}w^{-1})$, where $w$ is the long Weyl element, is self-adjoint (one can compute directly in explicit coordinates given in [20], Proposition 9.2.1 or Proposition 6.3.1). We can interpret the Fricke involution as
@@
where $m=A_{N}w^{-1}$, that is, as the composition of the dualising map above with the left-action of $m$. Since the measure on $\mathbb{H}^{n}$ is $\operatorname{GL}_{n}(\mathbb{R})$-invariant, we can make the same explicit computations and change of coordinates as for the dualising map. Since $A_{N}$ is diagonal, we easily deduce the conclusion $W_{N}^{\ast}=W_{N}$. Moreover, this interpretation of the Fricke involution and [4, (45)] also prove the statement about the spectral parameters of $\phi$.
To prove cuspidality it is best to work adelically, though this can be reduced again to noting the relation between $W_{N}$ and the dualising map. Namely, the form $W_{N}\phi$ generates the contragredient of the representation generated by $\phi$, which is known to be cuspidal (see e.g. [15, Prop. 3.3.4]). From this perspective, it is also easy to see that $W_{N}\phi$ is a newform. In the interest of brevity, we leave out the details of adelisation. [^4] Based on our definition, the adelic formulation of these operators and their application in computing epsilon factors in functional equations has been worked out by Dalal and Gerbelli-Gauthier [17, Sec. 3].
@@
As explained in [33], all automorphisms in the case $n>2$ are constructed out of inner automorphisms, radial automorphisms, and the inverse-transpose automorphism. Inner automorphisms cannot contribute, since we have proved that the normaliser of $\Gamma_{0}(N)$ is trivial; radial automorphisms are trivial in our context, since we consider only automorphic forms that are invariant under the centre of $\operatorname{GL}_{n}(\mathbb{R})$; and the inverse-transpose automorphism is precisely the basis for the definition given in this note.
4.2.1. Uniqueness of the Fricke involution
The theory of Atkin-Lehner operators for $\Gamma_{0}(N)$ shows some weaknesses already in the well-understood case $n=2$. Indeed, one can only define Atkin-Lehner operators for divisors $M$ of the level $N$, such that $M$ and $N/M$ are coprime. More precisely, there are no operators induced by matrices with determinant equal to $M\mid N$, such that $(M,N/M)\neq 1$ (see [2, p. 138]).
This phenomenon creates difficulties in applications when considering powerful levels, as already noted in the historical context of the sup-norm problem. Here, we see that these difficulties only get more problematic in higher rank. In fact, the only Atkin-Lehner operator for $n>2$, according to our definition, is the Fricke involution.
Corollary 4.4.
@@ 4.2.1. Uniqueness of the Fricke involution
This follows from Theorem 4 on the triviality of the normaliser of $\Gamma_{0}(N)$. Indeed, let $g$ be as in the statement and recall the definition of the matrix $A_{N}$ that gives the Fricke involution. We have that
$A_{N}\Gamma_{0}(N)^{T}A_{N}^{-1}=\Gamma_{0}(N),$
but also
$A_{N}\Gamma_{0}(N)^{T}A_{N}^{-1}=A_{N}g^{-1}\Gamma_{0}(N)gA_{N}^{-1}.$
Thus, $gA_{N}^{-1}$ belongs to the normaliser of $\Gamma_{0}(N)$ and the result follows. ∎